If y^-4 = 1/25 and y<0, then y^3 - y = ?
Answer by stanbon(75887)

If y^-4 = 1/25 and y<0, then y^3 - y = ?
------
(y^-4)^(-1/4) = (1/25)^(-1/4)
y = (1/5)^2 ^(-1/4) = (1/5)^-1/2 = (5)^(1/2) = sqrt(5)
-------------------------------
Ans: y^3 - y = 5^(3/2) - 5^(1/2)
----------
Cheers,
Stan H.
----------------